Hello. I would appreciate an explanation to the following question:
A playground is being designed where children can interact with their friends in certain combinations.
If there is 1 child, there can be 0 interactions.
If there are 2 children, there can be only 1 interaction.
If there are 3 children , there can be 5 interactions.
If there are 4 children, there can be 14 interactions.
Which recursive pattern represents the pattern?
asubn=asubn-1+2^(n-1)
asubn=asubn+(n-1)^2
asubn=asubn+2(n-1)
asubn=asubn+(2n-1)

So far I have figured out by setting up this basic pattern chart that:

n a
1 0 +1
2 1 +4
3 5 +9
4 14
Is the equation always the same? or does it change with the given information. I got this wrong and would appreciate understanding why? I understand how to make derive the sequence information. What if I need the 50th term? I can easily get the ninth term as in the previous question on this assignment. But having trouble understanding how to write the formula.
I appreciate you help. This is a wonderful resource for students who take math seriously because the tutors answer the questions so thoroughly. Again, thank you in advance.
Justina

Answer by Theo(3458)

the formula that works with the data is as follows:

a[n] = a[n-1] + (n-1)^2

here's a table of values based on that formula.
the table starts with a[1] being given as 0.

        n        (n-1)^2         a[n]

        1        n/a            given as 0
        2        1^2 = 1        1 + 0 = 1
        3        2^2 = 4        4 + 1 = 5
        4        3^2 = 9        9 + 5 = 14
        5        4^2 = 16       16 + 14 = 30
        6        5^2 = 25       25 + 30 = 55
        ...


you are given that the a[1] = 0
the rest of them are derived from the formula.
the same formula is used every time.
let me know if this solves the problem for you.